### v1

# assumption[theorem]Assumption

Abstract
We prove a Tauberian transfer principle showing that for any compact closed Riemannian manifold $(M^{d},g)$ the affine spectral encoding $C=\pi-\epsilon\lambda$ transports Laplacian Weyl asymptotics to a Weyl law in the edge variable in the bulk regime $C\to$-$\infty$ (equivalently, $(\pi-C)/\epsilon\to\infty$): $N_{\mu_{C}}(C)\sim\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2}$ and $\rho_{\mathrm{bulk}}(C)\sim\tfrac{d}{2}\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{(d-2)/2}$, so that $d$ and the Weyl constant $\gamma_{d}$ are recoverable from one–dimensional edge–variable data. Conversely, a bulk power law $N_{\mu_{C}}(C)\ \\ \sim\!A(\pi-C)^{\alpha}$ as $C\to$-$\infty$ implies $d=2\alpha$ and $\gamma_{d}=A\,\epsilon^{d/2}$. We establish the uniqueness of the affine rule among polynomial–type encodings $g(\lambda)=a-b\lambda^{k}L(\lambda)$ (the edge–variable exponent forces $k=1$) and stability under small perturbations $C=\pi-\epsilon\lambda+\delta(\lambda)$ with $\delta(\lambda)=o(\lambda)$. For constant–curvature model spaces we record strengthened correspondences for heat traces and spectral zeta, $H_{\mathrm{edge}}(s)=\Theta_{\Delta}(\epsilon s)$ and $\zeta_{\mathrm{edge}}(u)=\epsilon^{-u}\zeta_{\Delta}(u)$, and we realize multiplicities via generalized one-dimensional models (Krein strings). When a Weyl remainder $O(\Lambda^{(d-1)/2})$ is available, it transfers to a bulk remainder $O((\pi-C)^{(d-1)/2})$ in the CC–variable.
keywords:
weyl law; bulk-edge asymptotics; counting function; affine spectral encoding; laplace–beltrami operator; spectral zeta function; tauberian theorems; stability
1 Introduction
Theorem 1.1 (Main Theorem).
Let $(M^{d},g)$ be a smooth compact Riemannian manifold without boundary, and let $C=\pi-\epsilon\lambda$ be the affine encoding of the Laplace spectrum. Then:
(i) A Weyl law in the edge variable holds in the bulk regime $C\to$-$\infty$ (equivalently $(\pi-C)/\epsilon\to\infty$), yielding
$N_{\mu_{C}}(C)\sim\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2},\qquad\rho_{\mathrm{bulk}}(C)\sim\tfrac{d}{2}\,\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{(d-2)/2},$ (1.1)
so that $d$ and $\gamma_{d}$ are recoverable from one–dimensional edge–variable data.
(ii) Among polynomial-type encodings $g(\lambda)\sim a-b\lambda^{k}L(\lambda)$ with $L\in\mathrm{RV}_{0}$, only the asymptotically affine case $k=1$ preserves the geometric edge–variable density exponent (d2)/2(d-2)/2.
(iii) The edge–variable exponents are stable under perturbations $C=\pi-\epsilon\lambda+\delta(\lambda)$ with $\delta(\lambda)=o(\lambda)$.
Context.
The statement above rigorously realizes a principle informally articulated in our earlier work, where the edge–side spectral density was proposed as a dimensional signature (sec. VI.A in [1]). The present paper proves this mechanism via an affine pushforward of the Laplace spectrum and Tauberian tools, and further establishes the uniqueness of the affine encoding among polynomial–type rules and stability under $o(\lambda)$ perturbations.
Remark 1.2 (Terminology).
We reserve edge for the geometric endpoint $C\uparrow\pi$, while bulk refers to $C\to$-$\infty$. The asymptotics transferred from Weyl’s law occur in the bulk regime.
2 Preliminaries and standing assumptions
2.1 Geometry and Laplace spectrum
{assumption}
[Smoothness and spectrum] Throughout, $M^{d}$ is a smooth, connected, compact Riemannian manifold without boundary, and $\Delta_{g}$ denotes the nonnegative Laplace–Beltrami operator on $L^{2}(M)$. It is essentially self-adjoint on $C^{\infty}(M)$ and has purely discrete spectrum $0=\lambda_{0}\leq\lambda_{1}\leq\cdots\to\infty$, each eigenvalue with finite multiplicity $m_{n}$; [2, Ch. III], [3, §§2–3], [5, Ch. XIII].
2.2 Spectral measures and counting
Definition 2.1 (Spectral measure and counting function).
Let $(M^{d},g)$ be a compact Riemannian manifold without boundary, and let
$\Delta_{g}\phi_{n}=\lambda_{n}\phi_{n},\qquad\|\phi_{n}\|_{L^{2}(M)}=1,$ (2.1)
with eigenvalues $0=\lambda_{0}\leq\lambda_{1}\leq\cdots\to\infty$ and multiplicities $m_{n}$. The associated spectral measure is
$\mu_{\Delta}:=\sum_{n}m_{n}\,\delta_{\lambda_{n}},$ (2.2)
and its counting function is
$N_{\Delta}(\Lambda):=\mu_{\Delta}((-\infty,\Lambda])=\#\{n:\lambda_{n}\leq\Lambda\},$ (2.3)
where multiplicities are counted [4, Thm. VII.2].
Lemma 2.2 (Counting function properties).
$N_{\Delta}$ is nondecreasing,
right–continuous with left limits (càdlàg), and has jumps of size $m_{n}$ at the eigenvalues $\lambda_{n}$. Hence $N_{\Delta}\in BV_{\mathrm{loc}}([0,\infty))$ and is the Stieltjes distribution function of $\mu_{\Delta}$.
Proof 2.3.
Since $N_{\Delta}(\Lambda)=\mu_{\Delta}((-\infty,\Lambda])$ with a purely atomic measure, all stated properties are immediate.
2.3 Affine edge encoding and edge variable
Fix $\epsilon>0$ and define $C_{n}=\pi-\epsilon\lambda_{n}$, the pushforward measure $\mu_{C}:=\sum_{n}m_{n}\,\delta_{C_{n}}$, the edge variable $y:=\pi-C\downarrow 0$, and
$N_{\mu_{C}}(C):=\mu_{C}([C,\pi))=\#\{n:\ C_{n}\geq C\}.$ (2.4)
We write
$\rho_{\mathrm{edge}}(C):=-\,\frac{d}{dC}N_{\mu_{C}}(C)\ (\geq 0)$ (2.5)
in the Stieltjes/distributional sense (or after a standard smoothing). When emphasizing the bulk regime $C\to$-$\infty$, we use the synonymous notation
$\rho_{\mathrm{bulk}}(C):=-\,\frac{d}{dC}N_{\mu_{C}}(C),$ (2.6)
i.e. the same distributional derivative, with the name indicating the asymptotic regime.
Lemma 2.4 (Monotonicity of the encoding).
Fix $\epsilon>0$ and set $C=\pi-\epsilon\lambda$. Then the map $\lambda\mapsto C$ is strictly decreasing. Consequently,
$\lambda_{i}<\lambda_{j}\ \Longleftrightarrow\ C_{i}>C_{j},\qquad\lambda_{i}=\lambda_{j}\ \Longleftrightarrow\ C_{i}=C_{j},$ (2.7)
so the pushforward preserves multiplicities and order.
Proof 2.5.
Immediate from $dC/d\lambda=-\epsilon<0$.
Remark 2.6.
This strict monotonicity is the key reason why multiplicities and ordering transfer exactly to the edge-variable encoding.
2.4 Exact composition identity
Lemma 2.7 (Exact composition identity).
For every $C<\pi$ one has
$N_{\mu_{C}}(C)=\#\{n:\lambda_{n}\leq(\pi-C)/\epsilon\}=N_{\Delta}\!\big((\pi-C)/\epsilon\big).$ (2.8)
Proof 2.8.
Since $C=\pi-\epsilon\lambda$ is strictly decreasing, $\{n:C_{n}\geq C\}=\{n:\lambda_{n}\leq(\pi-C)/\epsilon\}$, with multiplicities.
2.5 Regular variation toolkit
A positive function $f$ is in $\mathrm{RV}_{\alpha}$ if $f(tu)/f(t)\to u^{\alpha}$ as $t\to\infty$ for each $u>0$. If $f(x)=x^{\alpha}\ell(x)$ with $\ell\in\mathrm{RV}_{0}$ (slowly varying), then $f\in\mathrm{RV}_{\alpha}$; [9, §§1.2–1.7].
2.6 Standing assumptions
{assumption}
[Standing framework] We assume throughout:
(i) $(M^{d},g)$ satisfies Assumption 2.1;
(ii) Weyl’s law holds
$N_{\Delta}(\Lambda)\sim\gamma_{d}\,\Lambda^{d/2}\qquad(\Lambda\to\infty),$ (2.9)
with the classical remainder $O(\Lambda^{(d-1)/2})$ when available; [6, 7] and, for refinements and boundary effects, [8, Ch. 1];
(iii) a fixed scaling parameter $\epsilon>0$ defines the affine encoding $C=\pi-\epsilon\lambda$;
(iv) we write $\omega_{d}:=\mathrm{Vol}(B_{d})=\pi^{d/2}/\Gamma(\tfrac{d}{2}+1)$ for the volume of the unit ball in $\mathbb{R}^{d}$;
(v) in Tauberian steps we tacitly use standard smoothings and the monotone density theorem for regularly varying functions [9, Thm. 1.7.2].
For manifolds with boundary we appeal to the two-term Weyl expansion; [8, Ch. 1], [7, Ch. 8].
3 Edge-Variable Weyl transfer: main results
3.1 Edge-variable Weyl (Abelian direction)
Theorem 3.1 (Edge-variable Weyl and reconstruction).
Let $(M^{d},g)$ be compact without boundary and suppose the Laplace–Beltrami counting function satisfies Weyl’s law
$N_{\Delta}(\Lambda)\sim\gamma_{d}\,\Lambda^{d/2}\qquad(\Lambda\to\infty),$ (3.1)
with $\gamma_{d}>0$. Fix $\epsilon>0$, encode $C=\pi-\epsilon\lambda$, and set
$N_{\mu_{C}}(C):=\#\{n:\ C_{n}\geq C\}\quad(C<\pi),\qquad\rho_{\mathrm{bulk}}(C):=-\frac{d}{dC}N_{\mu_{C}}(C)\ (\geq 0)$ (3.2)
in the Stieltjes/distributional sense (or after a standard smoothing). Then, as $C\to$-$\infty$ (equivalently, $(\pi-C)/\epsilon\to\infty$),
$N_{\mu_{C}}(C)\sim\gamma_{d}\,\epsilon^{-d/2}\,(\pi-C)^{d/2},$ (3.3)
and
$\rho_{\mathrm{bulk}}(C)\sim\frac{d}{2}\,\gamma_{d}\,\epsilon^{-d/2}\,(\pi-C)^{\frac{d-2}{2}}.$ (3.4)
Conversely, if for some $A>0$ and $\alpha>0$,
$N_{\mu_{C}}(C)\sim A\,(\pi-C)^{\alpha}\qquad(C\to-\infty),$ (3.5)
and $N_{\mu_{C}}$ is ultimately monotone, then necessarily
$d=2\alpha,\qquad\gamma_{d}=A\,\epsilon^{d/2}.$ (3.6)
Proof 3.2.
Since $C=\pi-\epsilon\lambda$ is strictly decreasing, the exact identity
$N_{\mu_{C}}(C)=N_{\Delta}\!\Big(\frac{\pi-C}{\epsilon}\Big)\qquad(C<\pi)$ (3.7)
holds with multiplicities (Lemma 2.7). With $\Lambda=(\pi-C)/\epsilon$ and Weyl’s law we get (3.3):
$N_{\mu_{C}}(C)=N_{\Delta}(\Lambda)\sim\gamma_{d}\,\Lambda^{d/2}=\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2}.$ (3.8)
For the density, put $y:=\pi-C\to\infty$ and $F(y):=N_{\mu_{C}}(\pi-y)=N_{\Delta}(y/\epsilon)$. Then
$F(y)\sim\gamma_{d}\,\epsilon^{-d/2}\,y^{d/2}=y^{d/2}\,\ell(y),\quad\text{with }\ \ell(y)\equiv\gamma_{d}\,\epsilon^{-d/2}\in\mathrm{RV}_{0}.$ (3.9)
By smoothing on scales $h=o(y)$ (Prop. 3.18) we may assume eventual monotonicity required by the monotone density theorem; the asymptotics are unchanged. By Karamata’s differentiation (monotone density) theorem for regularly varying functions (Bingham–Goldie–Teugels [9, §§1.4–1.7, Thm. 1.7.2]), if $F(y)=y^{\alpha}\ell(y)$ with $\alpha>0$ and $\ell\in\mathrm{RV}_{0}$ is ultimately monotone, then $F^{\prime}(y)\sim\alpha y^{\alpha-1}\ell(y)$. Hence
$F^{\prime}(y)\sim\frac{d}{2}\,\gamma_{d}\,\epsilon^{-d/2}\,y^{\frac{d-2}{2}}.$ (3.10)
Since $\rho_{\mathrm{bulk}}(C)=-\frac{d}{dC}N_{\mu_{C}}(C)=\frac{d}{dy}F(y)$ with $y=\pi-C$, (3.4) follows and $\rho_{\mathrm{bulk}}\geq 0$.
For the converse, with $\Lambda=(\pi-C)/\epsilon$,
$N_{\Delta}(\Lambda)=N_{\mu_{C}}(\pi-\epsilon\Lambda)\sim A(\epsilon\Lambda)^{\alpha}=A\,\epsilon^{\alpha}\,\Lambda^{\alpha}\qquad(\Lambda\to\infty),$ (3.11)
so necessarily $\alpha=d/2$ and $\gamma_{d}=A\,\epsilon^{d/2}$.
Corollary 3.3 (Practical recovery from bulk edge-variable data).
Let $y:=\pi-C\to\infty$. If
$N_{\mu_{C}}(\pi-y)\sim A\,y^{\alpha}\quad(y\to\infty)$ (3.12)
with ultimate monotonicity (or after smoothing as in Prop. 3.18), then
$d=2\alpha,\qquad\gamma_{d}=A\,\epsilon^{d/2}.$ (3.13)
Equivalently, the log–log slope $\alpha=\lim_{y\to\infty}\frac{d\log N_{\mu_{C}}(\pi-y)}{d\log y}$ yields $d=2\alpha$, and
$\gamma_{d}=\lim_{y\to\infty}\frac{N_{\mu_{C}}(\pi-y)}{y^{d/2}}\ \epsilon^{d/2}.$ (3.14)
The same holds if one uses the (smoothed/Stieltjes) bulk density $\rho_{\mathrm{bulk}}(C)\sim B\,y^{\alpha-1}$ with $y=\pi-C$, where $d=2\alpha$ and $B=\tfrac{d}{2}\gamma_{d}\,\epsilon^{-d/2}$.
3.2 Weyl remainder in the edge variable (bulk regime)
Proposition 3.4 (Transfer of remainder).
If $N_{\Delta}(\Lambda)=\gamma_{d}\Lambda^{d/2}+O(\Lambda^{(d-1)/2})$ as $\Lambda\to\infty$ (see [6, 7, 8]), then, as $C\to$-$\infty$ (equivalently, $\Lambda=(\pi-C)/\epsilon\to\infty$),
$N_{\mu_{C}}(C)=\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2}+O\!\left(\epsilon^{-(d-1)/2}(\pi-C)^{(d-1)/2}\right),$ (3.15)
with an implied constant depending only on $(M,g)$, hence uniform for $\epsilon$ in compact subsets of $(0,\infty)$.
Proof 3.5.
Apply Lemma 2.7 with $\Lambda=(\pi-C)/\epsilon$ and propagate the remainder. The substitution $\Lambda=(\pi-C)/\epsilon$ and monotonicity preserve both the order and form of the remainder.
Proposition 3.6 (Uniformity in $\epsilon$ in the bulk regime).
Fix $0<\epsilon_{0}<\epsilon_{1}<\infty$ and let $Y:=(\pi-C)/\epsilon$. Then, as $Y\to\infty$ (equivalently $C\to$-$\infty$),
$\sup_{\epsilon\in[\epsilon_{0},\epsilon_{1}]}\left|\frac{N_{\mu_{C}}(C)}{\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2}}-1\right|\;=\;\sup_{\epsilon\in[\epsilon_{0},\epsilon_{1}]}\left|\frac{N_{\Delta}(Y)}{\gamma_{d}\,Y^{d/2}}-1\right|\ \longrightarrow\ 0.$ (3.16)
Proof 3.7.
By Lemma 2.7, $N_{\mu_{C}}(C)=N_{\Delta}\!\big((\pi-C)/\epsilon\big)=N_{\Delta}(Y)$ with $Y=(\pi-C)/\epsilon$. Hence
$\sup_{\epsilon\in[\epsilon_{0},\epsilon_{1}]}\left|\frac{N_{\mu_{C}}(C)}{\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2}}-1\right|=\left|\frac{N_{\Delta}(Y)}{\gamma_{d}\,Y^{d/2}}-1\right|\xrightarrow[Y\to\infty]{}0,$ (3.17)
since $N_{\Delta}(Y)\sim\gamma_{d}Y^{d/2}$.
Proposition 3.8 (Remainder for bulk density).
If $N_{\Delta}(\Lambda)=\gamma_{d}\Lambda^{d/2}+O(\Lambda^{(d-1)/2})$ as $\Lambda\to\infty$, then after standard smoothing (as in Prop. 3.18), the bulk density satisfies
$\rho_{\mathrm{bulk}}(C)=\tfrac{d}{2}\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{(d-2)/2}+O\!\left(\epsilon^{-(d-2)/2}(\pi-C)^{(d-3)/2}\right),$ (3.18)
as $C\to$-$\infty$, using the monotone density theorem [9, Thm. 1.7.2].
Proof 3.9.
Apply the monotone density theorem to the remainder estimate from Proposition 3.4, using the same smoothing procedure as in Proposition 3.18.
3.3 Tauberian reconstruction (inverse direction)
We adopt the edge–counting convention
$N_{\mu_{C}}(C)\;:=\;\#\{\ell:\;C_{\ell}\geq C\}\qquad(C<\pi),$ (3.19)
which counts the atoms of $\mu_{C}$ to the right of $C$ (towards the edge $C=\pi$).
Lemma 3.10 (Change–of–variables dictionary).
For every $C<\pi$ one has the identity
$N_{\mu_{C}}(C)\;=\;\#\{\ell:\;C_{\ell}\in[C,\pi]\}\;=\;\#\{\ell:\;\lambda_{\ell}\leq\Lambda\}\;=\;N_{\Delta}(\Lambda),\qquad\Lambda=\frac{\pi-C}{\epsilon}.$ (3.20)
Proof 3.11.
Since $C_{\ell}=\pi-\epsilon\lambda_{\ell}$, the condition $C_{\ell}\in[C,\pi]$ is equivalent to $\pi-\epsilon\lambda_{\ell}\geq C$, i.e. $\lambda_{\ell}\leq(\pi-C)/\epsilon=\Lambda$. Counting with multiplicities gives the claim.
The lemma shows that the “bulk” regime $C\to$-$\infty$ is exactly the high–energy regime $\Lambda\to+\infty$. Hence Weyl’s law and the corresponding asymptotics for $N_{\mu_{C}}$ are equivalent.
Theorem 3.12 (Tauberian reconstruction of $d$ and $\gamma_{d}$).
Let $A>0$ and $\alpha>0$. The following are equivalent:
$\textup{(i)}\quad N_{\mu_{C}}(C)\ \sim\ A\,\epsilon^{-\alpha}\,(\pi-C)^{\alpha}\qquad(C\to-\infty),$
$\textup{(ii)}\quad N_{\Delta}(\Lambda)\ \sim\ A\,\Lambda^{\alpha}\qquad(\Lambda\to+\infty).$
In particular, under Weyl’s law $N_{\Delta}(\Lambda)\sim\gamma_{d}\,\Lambda^{d/2}$ one has
$N_{\mu_{C}}(C)\ \sim\ \gamma_{d}\,\epsilon^{-d/2}\,(\pi-C)^{d/2}\qquad(C\to-\infty),$ (3.21)
and conversely, observing the power–law exponent $\alpha$ and the coefficient $A$ of the normalized form $N_{\mu_{C}}(C)\sim A\,\epsilon^{-\alpha}(\pi-C)^{\alpha}$ determines
$d=2\alpha,\qquad\gamma_{d}=A.$ (3.22)
Proof 3.13.
By Lemma 3.10, $N_{\mu_{C}}(C)=N_{\Delta}\big((\pi-C)/\epsilon\big)$. Set $y:=\pi-C$ and $\Lambda:=y/\epsilon$. Then
$N_{\mu_{C}}(\pi-y)=N_{\Delta}(y/\epsilon).$ (3.23)
Thus $N_{\mu_{C}}(\pi-y)\sim A\,\epsilon^{-\alpha}y^{\alpha}$ as $y\to+\infty$ is equivalent to $N_{\Delta}(\Lambda)\sim A\,\Lambda^{\alpha}$ as $\Lambda\to+\infty$. The specialization with $\alpha=d/2$ and $A=\gamma_{d}$ is immediate.
Remark 3.14 (Practical estimators from bulk data).
Given samples of $N_{\mu_{C}}(C)$ for large negative $C$, the following consistent estimators follow from Theorem 3.12:
$\widehat{d}(C)\;=\;2\,\frac{d\log N_{\mu_{C}}(C)}{d\log(\pi-C)},\qquad\widehat{\gamma}_{d}(C)\;=\;\epsilon^{\widehat{d}(C)/2}\,\frac{N_{\mu_{C}}(C)}{(\pi-C)^{\widehat{d}(C)/2}}.$ (3.24)
In discrete data, replace the derivative by a local log–log slope. By Lemma 7.14, clustering contributes only bounded microscopic oscillations and does not affect the leading bulk asymptotics. See also Theorem 3.12 and Proposition 3.18.
Remark 3.15 (Edge vs. bulk).
The limit $C\to$-$\infty$ corresponds to $\Lambda\to+\infty$ (Weyl/high–energy, “bulk”), while the limit $C\uparrow\pi$ probes only finitely many low modes and is not governed by Weyl. To avoid confusion with the geometric “edge” $C\uparrow\pi$, we refer to $\rho_{\mathrm{bulk}}$ in the $C\to$-$\infty$ regime.
3.4 Boundary case: two-term bulk expansion
Proposition 3.16 (Two-term bulk expansion for manifolds with boundary).
Let $(M^{d},\partial M)$ be compact with smooth boundary and suppose
$N_{\Delta}(\Lambda)=\gamma_{d}\,\Lambda^{d/2}+\beta_{d}\,\Lambda^{(d-1)/2}+o\!\left(\Lambda^{(d-1)/2}\right)\qquad(\Lambda\to\infty).$ (3.25)
Here and below we consider the standard Dirichlet or Neumann boundary conditions; see [8, Ch. 1], [7, Ch. 8]. Then, as $C\to$-$\infty$ (equivalently, $y:=\pi-C\to\infty$),
$N_{\mu_{C}}(C)=\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2}+\beta_{d}\,\epsilon^{-(d-1)/2}(\pi-C)^{(d-1)/2}+o\!\left((\pi-C)^{(d-1)/2}\right).$ (3.26)
Proof 3.17.
Insert $\Lambda=(\pi-C)/\epsilon$ into the two-term Weyl expansion and use Lemma 2.7.
Proposition 3.18 (Stable smoothing of the bulk counting).
Let $\varphi\in C_{c}^{\infty}(\mathbb{R})$ be a nonnegative mollifier with $\int\varphi=1$, and set $\varphi_{h}(t):=h^{-1}\varphi(t/h)$. Define the smoothed counting function
$N_{\mu_{C}}^{(h)}(C):=(N_{\mu_{C}}*\varphi_{h})(C)=\int N_{\mu_{C}}(C-t)\,\varphi_{h}(t)\,dt.$ (3.27)
If $h=h(y)$ satisfies $h(y)=o(y)$ as $y=\pi-C\to\infty$, then
$N_{\mu_{C}}^{(h)}(C)\sim N_{\mu_{C}}(C)\sim\gamma_{d}\,\epsilon^{-d/2}\,y^{d/2},$ (3.28)
and the derivative $\frac{d}{dC}N_{\mu_{C}}^{(h)}(C)$ exists with
$\frac{d}{dC}N_{\mu_{C}}^{(h)}(C)=:\rho_{\mathrm{bulk}}^{(h)}(C)\ \sim\ \tfrac{d}{2}\,\gamma_{d}\,\epsilon^{-d/2}\,y^{\frac{d-2}{2}}.$ (3.29)
Proof 3.19.
Set $y:=\pi-C$ and $F(y):=N_{\mu_{C}}(\pi-y)$. By Lemma 2.7 and Weyl’s law we have
$F(y)=N_{\Delta}(y/\epsilon)\sim A\,y^{\alpha},\qquad A:=\gamma_{d}\,\epsilon^{-d/2},\ \ \alpha:=\tfrac{d}{2}>0,$ (3.30)
hence $F\in\mathrm{RV}_{\alpha}$ and is ultimately monotone. Let $\varphi\in C_{c}^{\infty}(\mathbb{R})$ be nonnegative with $\int\varphi=1$ and write $\varphi_{h}(t):=h^{-1}\varphi(t/h)$, where $h=h(y)$ satisfies $h(y)=o(y)$ as $y\to\infty$. Then
$N_{\mu_{C}}^{(h)}(C)=(N_{\mu_{C}}*\varphi_{h})(C)=(F*\varphi_{h})(y)=\int_{\mathbb{R}}F(y-t)\,\varphi_{h}(t)\,dt=\int_{\mathbb{R}}F(y-hu)\,\varphi(u)\,du,$ (3.31)
after the change of variables $t=hu$.
Since $\varphi$ has compact support, say $\mathrm{supp}\,\varphi\subset[-M,M]$, we have $|hu|\leq Mh=o(y)$. By the Uniform Convergence Theorem for regularly varying functions [9, Thm. 1.2.1] together with Potter bounds [9, Thm. 1.5.6], for every $\varepsilon>0$ there exists $y_{0}$ such that for all $y\geq y_{0}$ and $|u|\leq M$,
$\left|\frac{F(y-hu)}{F(y)}-1\right|\leq\varepsilon.$ (3.32)
Hence the integrand $\frac{F(y-hu)}{F(y)}\varphi(u)$ is dominated by an integrable bound and converges pointwise (in $u$) to $\varphi(u)$. By dominated convergence,
$\frac{N_{\mu_{C}}^{(h)}(C)}{F(y)}=\int\frac{F(y-hu)}{F(y)}\,\varphi(u)\,du\longrightarrow\int\varphi(u)\,du=1,$ (3.33)
so $N_{\mu_{C}}^{(h)}(C)\sim F(y)\sim\gamma_{d}\,\epsilon^{-d/2}\,y^{d/2}$ as $y\to\infty$.
For the derivative, convolution with a $C^{\infty}$ compactly supported kernel yields smoothness and
$\frac{d}{dC}N_{\mu_{C}}^{(h)}(C)=\frac{d}{dy}(F*\varphi_{h})(y)=(F^{\prime}*\varphi_{h})(y)=\int_{\mathbb{R}}F^{\prime}(y-hu)\,\varphi(u)\,du,$ (3.34)
where the derivative is in the (tempered) distribution sense. Since $F\in\mathrm{RV}_{\alpha}$ with $\alpha>0$ and is ultimately monotone, the monotone density theorem [9, Thm. 1.7.2] gives
$F^{\prime}(y)\sim\alpha A\,y^{\alpha-1}=\tfrac{d}{2}\,\gamma_{d}\,\epsilon^{-d/2}\,y^{\frac{d-2}{2}}\qquad(y\to\infty),$ (3.35)
so $F^{\prime}\in\mathrm{RV}_{\alpha-1}$ and is ultimately locally bounded and of one sign. Applying the same Uniform Convergence / dominated–convergence argument to $F^{\prime}$ yields
$\frac{d}{dy}(F*\varphi_{h})(y)=(F^{\prime}*\varphi_{h})(y)\sim F^{\prime}(y).$ (3.36)
Recalling $y=\pi-C$, we conclude
$\rho_{\mathrm{bulk}}^{(h)}(C)=-\frac{d}{dC}N_{\mu_{C}}^{(h)}(C)\sim\tfrac{d}{2}\,\gamma_{d}\,\epsilon^{-d/2}\,y^{\frac{d-2}{2}},$ (3.37)
as claimed.
4 Heat-trace and zeta correspondences in the edge variable
4.1 Heat trace transfer
Lemma 4.1 (Heat trace convergence and small-$t$ asymptotics).
On a smooth compact closed $(M^{d},g)$ the heat trace
$\Theta_{\Delta}(t):=\sum_{n}m_{n}e^{-t\lambda_{n}}$ (4.1)
converges for every $t>0$ and admits, as $t\downarrow 0$, the Seeley–DeWitt expansion
$\Theta_{\Delta}(t)\sim(4\pi t)^{-d/2}\sum_{j=0}^{\infty}a_{j}\,t^{j/2},$ (4.2)
where $a_{0}=\mathrm{Vol}(M)$ and $a_{j}$ are integrals of local curvature invariants; [3] Define the spectral zeta function by
$\zeta_{\Delta}(u):=\sum_{n}m_{n}\,\lambda_{n}^{-u}\qquad(\Re u>d/2).$ (4.3)
Consequently, $\zeta_{\Delta}$ admits a meromorphic continuation to $\mathbb{C}$ with at most simple poles located at $u=\frac{d-j}{2}$ for $j=0,1,2,\dots$.
Lemma 4.2 (Bounds for heat coefficients).
For each $j\geq 0$ there exists a universal polynomial $P_{j}$ in the curvature tensor and its covariant derivatives up to order $j$ such that
$a_{j}=\int_{M}P_{j}(\mathrm{Rm},\nabla\mathrm{Rm},\dots,\nabla^{j}\mathrm{Rm})\,d\mathrm{vol}_{g}.$ (4.4)
In particular, on a fixed compact $(M,g)$ one has $|a_{j}|\leq C_{j}(M,g)$.
Proof 4.3.
See, e.g., [10, Ch. 4] and [11].
Proposition 4.4 (Heat and zeta transfer).
Assume the purely affine encoding $C_{n}=\pi-\epsilon\lambda_{n}$ (i.e. $\delta\equiv 0$). With $\Theta_{\Delta}$ and $\zeta_{\Delta}$ as above, define
$H_{\mathrm{edge}}(s):=\sum_{n}m_{n}\,e^{-s(\pi-C_{n})},\qquad\zeta_{\mathrm{edge}}(u):=\sum_{n}m_{n}\,(\pi-C_{n})^{-u}.$ (4.5)
Then for all $s>0$ and $\Re u>d/2$,
$H_{\mathrm{edge}}(s)=\Theta_{\Delta}(\epsilon s),\qquad\zeta_{\mathrm{edge}}(u)=\epsilon^{-u}\,\zeta_{\Delta}(u).$ (4.6)
Hence the pole structure of $\zeta_{\mathrm{edge}}$ is identical to $\zeta_{\Delta}$.
Proof 4.5.
Since $\pi-C_{n}=\epsilon\lambda_{n}$ for all $n$, we have $H_{\mathrm{edge}}(s)=\sum_{n}m_{n}e^{-s(\pi-C_{n})}=\sum_{n}m_{n}e^{-\epsilon s\,\lambda_{n}}=\Theta_{\Delta}(\epsilon s)$, and $\zeta_{\mathrm{edge}}(u)=\sum_{n}m_{n}(\pi-C_{n})^{-u}=\epsilon^{-u}\sum_{n}m_{n}\lambda_{n}^{-u}=\epsilon^{-u}\zeta_{\Delta}(u)$.
Remark 4.6 (Effect of non-affine perturbations).
If $C_{n}=\pi-\epsilon\lambda_{n}+\delta(\lambda_{n})$ with $\delta(\lambda)=o(\lambda)$, then
$H_{\mathrm{edge}}(s)=\sum_{n}m_{n}e^{-\epsilon s\,\lambda_{n}}\,e^{\,s\,\delta(\lambda_{n})},\qquad\zeta_{\mathrm{edge}}(u)=\sum_{n}m_{n}\big(\epsilon\lambda_{n}-\delta(\lambda_{n})\big)^{-u}.$ (4.7)
Thus the exact identities of Proposition 4.4 are lost, although for asymptotic questions one recovers the leading behavior under the hypotheses of Section 6.
4.2 Spectral zeta transfer
Corollary 4.7.
For $\Re u>d/2$, $\zeta_{\mathrm{edge}}(u)=\epsilon^{-u}\,\zeta_{\Delta}(u)$. Consequently, the meromorphic continuation and residues of $\zeta_{\mathrm{edge}}$ are inherited from $\zeta_{\Delta}$ under the trivial scaling by $\epsilon^{-u}$.
Proof 4.8.
Immediate from Proposition 4.4.
5 Uniqueness of the affine encoding
5.1 Polynomial-type encodings and inverse asymptotics
We write $\rho_{\mu_{g}}(C):=-\frac{d}{dC}N_{\mu_{g}}(C)$ in the Stieltjes/distributional sense (or after a standard smoothing).
Theorem 5.1 (Uniqueness among polynomial-type encodings).
Let $g(\lambda)\sim a-b\,\lambda^{k}L(\lambda)$ as $\lambda\to\infty$ with $a\in\mathbb{R}$, $b>0$, $k>0$, and $L\in\mathrm{RV}_{0}$ (slowly varying). Assume $g$ is ultimately strictly decreasing so that $a$ is the upper edge of the pushforward measure $\mu_{g}$ of $\mu_{\Delta}$ under $g$ (by strict monotonicity of $g$; cf. Lemma 2.4 for the affine case). Put $\phi(\lambda):=a-g(\lambda)$ and $x:=a-C\downarrow 0$. Let $\Lambda$ be the asymptotic inverse of $\phi$. Then, as $C\uparrow a$,
$N_{\mu_{g}}(a-x)=N_{\Delta}(\Lambda(x))\in\mathrm{RV}_{\,d/(2k)},\qquad\rho_{\mu_{g}}(C)\in\mathrm{RV}_{\,d/(2k)-1}.$ (5.1)
In particular, to reproduce the edge-variable density exponent (d2)/2(d-2)/2 of (3.4) one must have $k=1$, i.e. $g(\lambda)=a-b\lambda(1+o(1))$.
Proof 5.2.
We have $\phi(\lambda)=a-g(\lambda)\sim b\,\lambda^{k}L(\lambda)$, so $\phi\in\mathrm{RV}_{k}$ and is ultimately increasing. By inversion theory for regularly varying functions (de Bruijn conjugacy; [9, §1.5, Thm. 1.5.12]), there exists an asymptotic inverse $\Lambda\in\mathrm{RV}_{1/k}$ with
$\Lambda(x)\sim\Big(\frac{x}{b}\Big)^{1/k}\,L^{\ast}(x)\qquad(x\downarrow 0),$ (5.2)
where $L^{\ast}\in\mathrm{RV}_{0}$ is characterized by
$\big(L^{\ast}(x)\big)^{k}\,L\!\Big(\Big(\frac{x}{b}\Big)^{\!1/k}L^{\ast}(x)\Big)\ \longrightarrow\ 1\qquad(x\downarrow 0).$ (5.3)
Monotonicity yields the exact composition $N_{\mu_{g}}(a-x)=N_{\Delta}(\Lambda(x))$. Applying Weyl’s law gives
$N_{\mu_{g}}(a-x)\sim\gamma_{d}\,\Lambda(x)^{d/2}\in\mathrm{RV}_{\,d/(2k)}\qquad(x\downarrow 0).$ (5.4)
Let $F(x):=N_{\mu_{g}}(a-x)$. By the monotone density theorem ([9, §1.7, Thm. 1.7.2]), $F^{\prime}(x)\in\mathrm{RV}_{\,d/(2k)-1}$. Since $\rho_{\mu_{g}}(C)=-\frac{d}{dC}N_{\mu_{g}}(C)=\frac{d}{dx}F(x)$ with x=aCx=a-C, the edge density exponent is d/(2k)1d/(2k)-1. Matching this with (d2)/2(d-2)/2 forces d/(2k)1=(d2)/2d/(2k)-1=(d-2)/2, i.e. $k=1$. Slowly varying factors $L,L^{\ast}$ affect the prefactor but not the exponent.
Corollary 5.3 (Nonlinear distortion of the edge-variable exponent).
Under the assumptions of Theorem 5.1, if $k\neq 1$ then the edge-variable density exponent equals $d/(2k)-1\neq(d-2)/2$. For instance, for $k=2$ one has
$\rho_{\mu_{g}}(C)\in\mathrm{RV}_{\,d/4-1},$ (5.5)
which mismatches the geometric edge-variable exponent for any $d\geq 1$.
Corollary 5.4 (Affine optimality for the edge-variable exponent).
Among encodings $g(\lambda)=a-b\lambda^{k}L(\lambda)$ with $b>0$, $k>0$, $L\in\mathrm{RV}_{0}$, the only ones reproducing both the correct edge-variable exponent (d2)/2(d-2)/2 and a finite, nonzero edge-variable prefactor are those with $k=1$. If moreover $L(\lambda)\to 1$, the prefactor matches $\tfrac{d}{2}\gamma_{d}\,b^{-d/2}$; any $L\not\to 1$ induces a slowly varying multiplicative distortion.
Proof $5.$
Immediate from Theorem 5.1 and the monotone density theorem [9, §1.7].
5.2 Edge exponents and the necessity of $k=1$
Proposition 5.6 (Edge-variable exponents under $g(\lambda)\sim a-b\lambda^{k}L(\lambda)$).
Let $g(\lambda)\sim a-b\,\lambda^{k}L(\lambda)$ as $\lambda\to\infty$ with $a\in\mathbb{R}$, $b>0$, $k>0$, and $L\in\mathrm{RV}_{0}$, and assume $g$ is ultimately strictly decreasing (by strict monotonicity of $g$; cf. Lemma 2.4 for the affine case). Put $x:=a-C\downarrow 0$ and let $\mu_{g}$ be the pushforward of $\mu_{\Delta}$ by $g$. Then
$N_{\mu_{g}}(a-x)\in\mathrm{RV}_{\,d/(2k)},\qquad\rho_{\mu_{g}}(a-x)\in\mathrm{RV}_{\,d/(2k)-1}.$ (5.6)
In particular, the edge-variable counting exponent equals $d/(2k)$ and the edge-variable density exponent equals d/(2k)1d/(2k)-1.
Proof 5.7.
This is an immediate consequence of Theorem 5.1: writing $\phi(\lambda)=a-g(\lambda)\in\mathrm{RV}_{k}$ and letting $\Lambda$ be an asymptotic inverse, one has $N_{\mu_{g}}(a-x)=N_{\Delta}(\Lambda(x))\sim\gamma_{d}\,\Lambda(x)^{d/2}\in\mathrm{RV}_{d/(2k)}$, and then the monotone density theorem yields the claim for $\rho_{\mu_{g}}$.
Corollary 5.8 (Necessity of $k=1$ to match the geometric edge-variable exponent).
The geometric edge-variable density exponent is (d2)/2(d-2)/2. Matching this with the value d/(2k)1d/(2k)-1 from Proposition 5.6 forces $k=1$. Hence any polynomial-type encoding preserving the edge-variable exponent must be asymptotically affine:
$g(\lambda)=a-b\lambda\,L(\lambda),\qquad L\in\mathrm{RV}_{0}.$ (5.7)
Slow variation $L$ cannot alter the exponent (only the slowly varying prefactor).
Remark 5.9 (What happens for $k\neq 1$).
If $k>1$ the edge-variable density exponent d/(2k)1d/(2k)-1 is strictly smaller than (d2)/2(d-2)/2, yielding an artificially thinner edge density; if $k<1$ it is strictly larger, yielding an artificially thicker edge density. In both cases the mismatch cannot be corrected by any slowly varying factor $L$.
Corollary 5.10 (Estimating $k$ from edge-variable data).
Suppose near the upper edge $C\uparrow a$ the counting behaves as
$N_{\mu_{g}}(C)\sim A\,(a-C)^{\alpha}\quad\text{with }\alpha>0,$ (5.8)
or, equivalently, the edge-variable density behaves as
$\rho_{\mu_{g}}(C)\sim B\,(a-C)^{\beta}\quad\text{with }\beta>-1.$ (5.9)
Then $\alpha=d/(2k)$ and $\beta=d/(2k)-1$, so
$k=\frac{d}{2\alpha}=\frac{d}{2(\beta+1)}.$ (5.10)
Thus the edge-variable log–log slope (of either $N$ or $\rho$) recovers $k$; matching the geometric value requires $k=1$.
5.3 Slow variation and prefactors
Proposition 5.11 (Slow variation does not change edge-variable exponents).
Let $g(\lambda)\sim a-b\,\lambda^{k}L(\lambda)$ with $b>0$, $k>0$, and $L\in\mathrm{RV}_{0}$, and assume $g$ is ultimately strictly decreasing. Put $\phi(\lambda):=a-g(\lambda)\sim b\lambda^{k}L(\lambda)$ and let $\Lambda$ be an asymptotic inverse of $\phi$. Then, with x:=aCx:=a-C,
$N_{\mu_{g}}(a-x)=N_{\Delta}(\Lambda(x))\in\mathrm{RV}_{\,d/(2k)},\qquad\rho_{\mu_{g}}(a-x)\in\mathrm{RV}_{\,d/(2k)-1}.$ (5.11)
In particular, $L\in\mathrm{RV}_{0}$ can modify only the prefactor by a slowly varying multiplicative factor; the exponents $d/(2k)$ and d/(2k)1d/(2k)-1 are unaffected.
Proof 5.12.
By de Bruijn inversion [9, Thm. 1.5.12], there exists $L^{\ast}\in\mathrm{RV}_{0}$ such that
$\Lambda(x)\sim\Big(\frac{x}{b}\Big)^{1/k}L^{\ast}(x)\quad\text{with}\quad\big(L^{\ast}(x)\big)^{k}\,L\!\Big(\Big(\tfrac{x}{b}\Big)^{1/k}L^{\ast}(x)\Big)\to 1.$ (5.12)
Then Weyl’s law gives
$N_{\mu_{g}}(a-x)\sim\gamma_{d}\,\Lambda(x)^{d/2}\sim\gamma_{d}\,b^{-d/(2k)}\,x^{d/(2k)}\,(L^{\ast}(x))^{d/2}\in\mathrm{RV}_{\,d/(2k)},$ (5.13)
and the monotone density theorem yields the claim for $\rho_{\mu_{g}}$.
Corollary 5.13 (Explicit prefactor with de Bruijn conjugate).
Under the assumptions of Proposition 5.11,
$N_{\mu_{g}}(a-x)$ $\sim\gamma_{d}\,b^{-d/(2k)}\,x^{d/(2k)}\,(L^{\ast}(x))^{d/2},$ (5.14)
$\rho_{\mu_{g}}(a-x)$ $\sim\tfrac{d}{2k}\,\gamma_{d}\,b^{-d/(2k)}\,x^{\frac{d}{2k}-1}\,(L^{\ast}(x))^{d/2},$ (5.15)
where $L^{\ast}\in\mathrm{RV}_{0}$ is the de Bruijn conjugate of $L$ associated to $\phi(\lambda)=b\lambda^{k}L(\lambda)$.
Remark 5.14 (Specializations for $k=1$).
When $k=1$ (the asymptotically affine case required by Corollary 5.8):
  1. 1.
    If $L(\lambda)\to\ell_{\infty}\in(0,\infty)$, then $L^{\ast}(x)\to\ell_{\infty}^{-1}$ and
    $N_{\mu_{g}}(a-x)$ $\sim\gamma_{d}\,(b\ell_{\infty})^{-d/2}\,x^{d/2},$ (5.16)
    $\rho_{\mu_{g}}(a-x)$ $\sim\tfrac{d}{2}\,\gamma_{d}\,(b\ell_{\infty})^{-d/2}\,x^{\frac{d-2}{2}}.$ (5.17)
    Thus a constant slow factor merely rescales the prefactor by $(\ell_{\infty})^{-d/2}$.
  2. 2.
    If $L(\lambda)=(\log\lambda)^{\alpha}(\log\log\lambda)^{\beta}$ for large $\lambda$, then $L^{\ast}(x)$ is again slowly varying and behaves (in the corresponding asymptotic regime) like $(\log x)^{-\alpha}(\log\log x)^{-\beta}$; hence $N_{\mu_{g}}$ and $\rho_{\mu_{g}}$ pick up a multiplicative modulation $(L^{\ast}(x))^{d/2}$ while their power exponents stay unchanged.
Corollary 5.15 (Affine normalization and matching of prefactors).
In the affine case $g(\lambda)=a-\epsilon\,\lambda\,L(\lambda)$ with $L\to 1$, the de Bruijn factor satisfies $L^{\ast}(x)\to 1$ and the edge-variable prefactors reduce to the canonical values
$N_{\mu_{g}}(a-x)$ $\sim\gamma_{d}\,\epsilon^{-d/2}\,x^{d/2},$ (5.18)
$\rho_{\mu_{g}}(a-x)$ $\sim\tfrac{d}{2}\,\gamma_{d}\,\epsilon^{-d/2}\,x^{\frac{d-2}{2}}.$ (5.19)
More generally, any $L\in\mathrm{RV}_{0}$ only changes these prefactors by the slowly varying factor $(L^{\ast}(x))^{d/2}$.
5.4 Uniqueness of the affine scaling from the bulk exponent
Let $g(\lambda)=a-\phi(\lambda)$ with $\phi(\lambda)>0$ ultimately increasing and $\phi(\lambda)\to\infty$. Write $x:=a-C\downarrow 0$ and let $\mu_{g}$ be the pushforward of $\mu_{\Delta}$ by $g$.
Theorem 5.16 (Bulk exponent singles out index $1$ in full $O$-regular variation).
Assume $\phi$ is $O$-regularly varying at $\infty$ with finite Matuszewska indices $0<\alpha_{\phi}^{-}\leq\alpha_{\phi}^{+}<\infty$ (Bingham–Goldie–Teugels [9, §2.1]), and is ultimately increasing. Let $\Lambda$ be an asymptotic inverse of $\phi$ near 0. Then:
  1. (i)
    $\Lambda$ is $O$-regularly varying at $0+$ with indices $1/\alpha_{\phi}^{+}\leq\mathrm{ind}(\Lambda)\leq 1/\alpha_{\phi}^{-}$.
  2. (ii)
    Consequently,
    $N_{\mu_{g}}(a-x)$ $=N_{\Delta}(\Lambda(x))\in O\text{-}\mathrm{RV}_{\,\left[d/(2\alpha_{\phi}^{+}),\,d/(2\alpha_{\phi}^{-})\right]},$ (5.20)
    $\rho_{\mu_{g}}(a-x)$ $\in O\text{-}\mathrm{RV}_{\,\left[d/(2\alpha_{\phi}^{+})-1,\,d/(2\alpha_{\phi}^{-})-1\right]}.$ (5.21)
  3. (iii)
    If, in fact, a genuine power law holds at the edge,
    $N_{\mu_{g}}(a-x)\sim A\,x^{d/2}\quad\text{or}\quad\rho_{\mu_{g}}(a-x)\sim B\,x^{(d-2)/2},$ (5.22)
    then necessarily $\alpha_{\phi}^{-}=\alpha_{\phi}^{+}=1$. Hence $\phi\in\mathrm{RV}_{1}$ and
    $\phi(\lambda)=b\,\lambda\,L(\lambda),\qquad b>0,\quad L\in\mathrm{RV}_{0},$ (5.23)
    so $g(\lambda)=a-b\lambda L(\lambda)$ is asymptotically affine (up to slow variation). In particular, the case $k=1$ of Theorem 5.1 is necessary and sufficient at the level of one-sided bulk power exponents.
Proof 5.17.
(i)–(ii) follow from standard inversion properties of $O$-regularly varying functions and bounding the inverse via the upper/lower Matuszewska indices [9, §§2.1–2.4]. Composing with Weyl’s law $N_{\Delta}(\Lambda)\sim\gamma_{d}\,\Lambda^{d/2}$ transfers the index interval. (iii) A single sharp power exponent for $N_{\mu_{g}}$ (or for $\rho_{\mu_{g}}$ via the monotone density theorem) forces the upper and lower indices of $\Lambda$ to coincide, hence those of $\phi$ coincide and equal $1$, i.e. $\phi\in\mathrm{RV}_{1}$. Then $\phi(\lambda)=b\lambda L(\lambda)$ with $L\in\mathrm{RV}_{0}$ by Karamata’s representation [9, Thm. 1.3.1], which yields the stated form of $g$.
Remark 5.18 (Rapid/subliminal distortions cannot reproduce the bulk power).
If $\phi$ is faster than any power (e.g. $\phi(\lambda)=e^{\lambda}$), then $\Lambda(x)\asymp\log(1/x)$ and $N_{\mu_{g}}(a-x)\asymp(\log(1/x))^{d/2}$, $\rho_{\mu_{g}}(a-x)\asymp x^{-1}(\log(1/x))^{d/2-1}$, which is not a power law. If $\phi(\lambda)=\lambda^{k}L(\lambda)$ with $k\neq 1$, Proposition 5.6 shows the exponent mismatch $d/(2k)-1\neq(d-2)/2$.
6 Stability under perturbations
6.1 Setup and inversion error
Lemma 6.1 (Explicit inversion error under tame perturbations).
Assume $C(\lambda)=\pi-\epsilon\lambda+\delta(\lambda)$ is eventually strictly decreasing, with
$|\delta(\lambda)|\leq\eta(\lambda)\,\lambda,\qquad\eta\in\mathrm{RV}_{0},\ \eta(\lambda)\to 0.$ (6.1)
Let $y:=\pi-C$ and denote by $\Lambda(C)$ the inverse of $\lambda\mapsto C(\lambda)$ for large $\lambda$ (equivalently, for large $y$). Then, as $C\to$-$\infty$ (i.e. $y\to+\infty$),
$\Lambda(C)=\frac{y}{\epsilon}\Big(1+O\big(\eta(y/\epsilon)\big)\Big).$ (6.2)
Consequently,
$N_{\mu_{C}}(C)=N_{\Delta}(\Lambda(C))=\gamma_{d}\,\epsilon^{-d/2}\,y^{d/2}\Big(1+O\big(\eta(y/\epsilon)\big)\Big),$ (6.3)
and the same relative error carries to $\rho_{\mathrm{bulk}}(C)$ after (Stieltjes) differentiation.
Proof 6.2.
Set $y:=\pi-C$ and fix $\Lambda_{0}$ so large that for all $\lambda\geq\Lambda_{0}$ one has $\eta(\lambda)\leq\epsilon/2$ and $C(\lambda)$ is strictly decreasing; consequently the inverse $\Lambda(C)$ is well defined for all $y\geq Y_{0}:=\pi-C(\Lambda_{0})$.
Step 1: two–sided inversion bounds. From $y=\epsilon\lambda-\delta(\lambda)$ and $|\delta(\lambda)|\leq\eta(\lambda)\lambda$ we get
$\lambda\big(\epsilon-\eta(\lambda)\big)\ \leq\ y\ \leq\ \lambda\big(\epsilon+\eta(\lambda)\big),$ (6.4)
hence, for all large $y$ (so that $\eta(\lambda)<\epsilon/2$),
$\frac{y}{\epsilon+\eta(\lambda)}\ \leq\ \lambda\ \leq\ \frac{y}{\epsilon-\eta(\lambda)}.$ (6.5)
Dividing by $y/\epsilon$ and using $(1\pm t)^{-1}=1\mp t+O(t^{2})$ yields
$\frac{\lambda}{y/\epsilon}\ \in\ \left[\,\frac{\epsilon}{\epsilon+\eta(\lambda)}\,,\,\frac{\epsilon}{\epsilon-\eta(\lambda)}\,\right]=\left[\,1-\frac{\eta(\lambda)}{\epsilon}+O\!\big(\eta(\lambda)^{2}\big),\ 1+\frac{\eta(\lambda)}{\epsilon}+O\!\big(\eta(\lambda)^{2}\big)\right].$ (6.6)
Therefore
$\lambda=\frac{y}{\epsilon}\Big(1+O\big(\eta(\lambda)\big)\Big)\qquad(y\to\infty),$ (6.7)
and in particular $\lambda/(y/\epsilon)\to 1$ as $y\to\infty$.
Step 2: replacing $\eta(\lambda)$ by $\eta(y/\epsilon)$. Since $\eta\in\mathrm{RV}_{0}$ (slowly varying), the Uniform Convergence Theorem (Bingham–Goldie–Teugels [9, Thm. 1.2.1]) gives
$\frac{\eta(\lambda)}{\eta(y/\epsilon)}\ \longrightarrow\ 1\qquad\text{whenever}\quad\frac{\lambda}{y/\epsilon}\ \longrightarrow\ 1.$ (6.8)
Using (6.7), this applies here, so $\eta(\lambda)=\eta(y/\epsilon)\,(1+o(1))$. Thus
$\Lambda(C)=\lambda=\frac{y}{\epsilon}\Big(1+O\big(\eta(y/\epsilon)\big)\Big)\qquad(y\to\infty),$ (6.9)
which is the first claimed estimate.
Step 3: asymptotics for $N_{\mu_{C}}$. By monotonicity of $C(\cdot)$ we have the exact identity $N_{\mu_{C}}(C)=\#\{\lambda_{\ell}\leq\Lambda(C)\}=N_{\Delta}(\Lambda(C))$. (Cf. Lemma 2.4.) Weyl’s law yields
$N_{\Delta}(\Lambda)=\gamma_{d}\,\Lambda^{d/2}\,(1+o(1))\qquad(\Lambda\to\infty).$ (6.10)
Insert (6.9): write $\Lambda=(y/\epsilon)\,(1+\theta(y))$ with $\theta(y)=O(\eta(y/\epsilon))$. Then
$\Lambda^{d/2}=\Big(\tfrac{y}{\epsilon}\Big)^{\!d/2}\,(1+\theta(y))^{d/2}=\Big(\tfrac{y}{\epsilon}\Big)^{\!d/2}\Big(1+O\big(\eta(y/\epsilon)\big)\Big),$ (6.11)
and hence
$N_{\mu_{C}}(C)=\gamma_{d}\,\epsilon^{-d/2}\,y^{d/2}\Big(1+O\big(\eta(y/\epsilon)\big)\Big)\qquad(y\to\infty).$ (6.12)
Step 4: bulk density. Define $F(y):=N_{\mu_{C}}(\pi-y)$, so $F$ is nondecreasing and $F(y)\sim A\,y^{\alpha}L(y)$ with
$A:=\gamma_{d}\,\epsilon^{-d/2},\quad\alpha:=\tfrac{d}{2}>0,\quad L(y):=1+O\big(\eta(y/\epsilon)\big).$ (6.13)
Because $\eta\in\mathrm{RV}_{0}$, the factor $L$ is slowly varying (and $L(y)\to 1$). By the monotone density (Karamata differentiation) theorem for regularly varying functions [9, §§1.4–1.7, Thm. 1.7.2], one has
$F^{\prime}(y)\ \sim\ \alpha\,A\,y^{\alpha-1}\,L(y)=\tfrac{d}{2}\,\gamma_{d}\,\epsilon^{-d/2}\,y^{\frac{d-2}{2}}\Big(1+O\big(\eta(y/\epsilon)\big)\Big),$ (6.14)
where the derivative is taken in the (Stieltjes/distributional) sense; equivalently, the same asymptotic holds for any standard smooth regularization of the density. Since $\rho_{\mathrm{bulk}}(C)=-\frac{d}{dC}N_{\mu_{C}}(C)=\frac{d}{dy}F(y)$ with $y=\pi-C$, this gives the stated estimate for $\rho_{\mathrm{bulk}}(C)$.
Combining the four steps completes the proof.
6.2 Preservation of edge-variable exponents
Proposition 6.3 (Stability under small perturbations: bulk formulation).
Let
$C(\lambda)=\pi-\epsilon\lambda+\delta(\lambda),\qquad\epsilon>0,$ (6.15)
where $\delta(\lambda)=\lambda\,\eta(\lambda)$ with $\eta\in\mathrm{RV}_{0}$ and $\eta(\lambda)\to 0$ as $\lambda\to\infty$. Assume moreover that $\delta$ is $C^{1}$ eventually, $\delta^{\prime}\in\mathrm{RV}_{0}$, and $\eta$ is ultimately monotone (equivalently, of bounded variation on a tail). Set $y:=\pi-C$ and let $\Lambda(C)$ be the inverse of $\lambda\mapsto C(\lambda)$. Then, as $C\to$-$\infty$ (equivalently $y\to+\infty$),
$\Lambda(C)$ $=\frac{y}{\epsilon}\Big(1+O\!\big(\eta(y/\epsilon)\big)\Big),$ (6.16)
$N_{\mu_{C}}(C)$ $=\gamma_{d}\,\epsilon^{-d/2}\,y^{d/2}\Big(1+O\!\big(\eta(y/\epsilon)\big)\Big),$ (6.17)
$\rho_{\mathrm{bulk}}(C)$ $=\tfrac{d}{2}\,\gamma_{d}\,\epsilon^{-d/2}\,y^{\frac{d-2}{2}}\Big(1+O\!\big(\eta(y/\epsilon)\big)\Big).$ (6.18)
hence the edge-variable (bulk) exponents are preserved and the prefactors are perturbed by a multiplicative $(1+O(\eta(y/\epsilon)))$ factor (in particular, by $(1+o(1))$). The $O(\cdot)$ terms are uniform for $\epsilon\in[\epsilon_{0},\epsilon_{1}]\subset(0,\infty)$.
Proof 6.4.
Write $y=\pi-C(\lambda)=\epsilon\lambda-\delta(\lambda)=\epsilon\lambda\,[1-\eta(\lambda)/\epsilon]$. Thus
$\lambda=\frac{y}{\epsilon}\,\Big[1-\eta(\lambda)/\epsilon\Big]^{-1}=\frac{y}{\epsilon}\,\Big(1+\tfrac{1}{\epsilon}\eta(\lambda)+O(\eta(\lambda)^{2})\Big).$ (6.19)
By the uniform convergence theorem for slowly varying functions [9, §1.8], replace $\eta(\lambda)$ by $\eta(y/\epsilon)$ when $\lambda\sim y/\epsilon$, yielding
$\Lambda(C)=\frac{y}{\epsilon}\Big(1+O(\eta(y/\epsilon))\Big),\qquad y\to\infty.$ (6.20)
Now apply Weyl’s law $N_{\Delta}(\Lambda)\sim\gamma_{d}\,\Lambda^{d/2}$ (valid as $\Lambda\to\infty$) and stability of regular variation under composition [9, §1.7] to get
$N_{\mu_{C}}(C)=N_{\Delta}(\Lambda(C))=\gamma_{d}\,\epsilon^{-d/2}\,y^{d/2}\Big(1+O(\eta(y/\epsilon))\Big).$ (6.21)
Set $F(y):=N_{\mu_{C}}(\pi-y)$. Then $F\in\mathrm{RV}_{d/2}$ and is ultimately monotone; by Karamata’s monotone density theorem [9, Thm. 1.7.2],
$F^{\prime}(y)\sim\tfrac{d}{2}\,\gamma_{d}\,\epsilon^{-d/2}\,y^{\frac{d-2}{2}}\Big(1+O(\eta(y/\epsilon))\Big).$ (6.22)
Since $\rho_{\mathrm{bulk}}(C)=-\frac{d}{dC}N_{\mu_{C}}(C)=\frac{d}{dy}F(y)$ with $y=\pi-C$, the stated asymptotics for $\rho_{\mathrm{bulk}}$ follow.
Lemma 6.5 (A convenient growth condition on $\delta$).
Assume $\delta(\lambda)=o(\lambda)$, $\delta$ is ultimately monotone, and there exists a slowly varying $\ell\in\mathrm{RV}_{0}$ with
$|\delta^{\prime}(\lambda)|\leq\frac{\ell(\lambda)}{\lambda}\quad\text{for large }\lambda.$ (6.23)
Then the conclusions of Lemma 6.1 and Proposition 6.3 hold (in the bulk formulation above) with $\eta=\ell$.
Proof $6.$
Integrating $|\delta^{\prime}(\lambda)|\leq\ell(\lambda)/\lambda$ yields $\delta(\lambda)=o(\lambda)$ with explicit control; then apply Lemma 6.1 and the proof of Proposition 6.3, using the uniform convergence theorem for slowly varying functions [9, §1.8].
6.3 Examples of admissible perturbations
Proposition 6.7 (Admissible perturbations and bulk error profiles).
Let $C(\lambda)=\pi-\epsilon\lambda+\delta(\lambda)$ be ultimately strictly decreasing and put $y:=\pi-C\to\infty$. The following choices of $\delta$ are admissible:
  1. (a)
    Logarithmic slope distortions. For $\alpha>0$ set
    $\delta(\lambda)=\frac{\lambda}{(\log(e\lambda))^{\alpha}},\qquad\eta(\lambda)=\frac{1}{(\log(e\lambda))^{\alpha}}\in\mathrm{RV}_{0}.$ (6.24)
    Then
    $\Lambda(C)=\frac{y}{\epsilon}\Big(1+O((\log(y/\epsilon))^{-\alpha})\Big),$ (6.25)
    $N_{\mu_{C}}(C)$ $=\gamma_{d}\epsilon^{-d/2}y^{d/2}\Big(1+O((\log(y/\epsilon))^{-\alpha})\Big),$ (6.26)
    $\rho_{\mathrm{bulk}}(C)$ $=\tfrac{d}{2}\gamma_{d}\epsilon^{-d/2}y^{\frac{d-2}{2}}\Big(1+O((\log(y/\epsilon))^{-\alpha})\Big).$ (6.27)
  2. (b)
    Iterated logarithms. For $\alpha>0$, $\beta\in\mathbb{R}$ let
    $\delta(\lambda)=\frac{\lambda}{(\log(e\lambda))^{\alpha}(\log\log(e^{e}\lambda))^{\beta}},\quad\eta(\lambda)=\frac{1}{(\log(e\lambda))^{\alpha}(\log\log(e^{e}\lambda))^{\beta}}\in\mathrm{RV}_{0}.$ (6.28)
    Then all conclusions in (a) hold with
    $O(\eta(y/\epsilon))=O\!\Big((\log(y/\epsilon))^{-\alpha}(\log\log(y/\epsilon))^{-\beta}\Big).$ (6.29)
  3. (c)
    General slowly varying multiplicative distortions. Let $L\in\mathrm{RV}_{0}$ be ultimately monotone with $L(\lambda)\to 0$, and put $\delta(\lambda)=\lambda L(\lambda)$ (so $\eta=L$). Then
    $\Lambda(C)$ $=\frac{y}{\epsilon}\big(1+O(L(y/\epsilon))\big),$ (6.30)
    $N_{\mu_{C}}(C)$ $=\gamma_{d}\epsilon^{-d/2}y^{d/2}\big(1+O(L(y/\epsilon))\big),$ (6.31)
    and the same relative error holds for $\rho_{\mathrm{bulk}}(C)$.
  4. (d)
    Bounded additive offsets. If $\delta(\lambda)=O(1)$ and $C$ is ultimately strictly decreasing, then
    $\Lambda(C)$ $=\frac{y}{\epsilon}+O(1),$ (6.32)
    $N_{\mu_{C}}(C)$ $=\gamma_{d}\epsilon^{-d/2}y^{d/2}\Big(1+O(y^{-1})\Big),$ (6.33)
    $\rho_{\mathrm{bulk}}(C)=\tfrac{d}{2}\gamma_{d}\epsilon^{-d/2}y^{\frac{d-2}{2}}\Big(1+O(y^{-1})\Big).$ (6.34)
  5. (e)
    Sublogarithmic additive terms. For $0<\beta<1$, let
    $\delta(\lambda)=(\log(e\lambda))^{\beta}.$ (6.35)
    Then $\delta(\lambda)=o(\lambda)$ and $\delta^{\prime}(\lambda)=\beta(\log(e\lambda))^{\beta-1}\lambda^{-1}\leq\ell(\lambda)/\lambda$ with $\ell(\lambda)=\beta(\log(e\lambda))^{\beta-1}\in\mathrm{RV}_{0}$, $\ell(\lambda)\to 0$. Hence
    $\Lambda(C)=\frac{y}{\epsilon}\Big(1+O((\log(y/\epsilon))^{\beta-1})\Big),$ (6.36)
    and the corresponding $N_{\mu_{C}}(C)$ and $\rho_{\mathrm{bulk}}(C)$ have the same relative error $O((\log(y/\epsilon))^{\beta-1})$.
  6. (f)
    Mild oscillatory but BV distortions. Let $L\in\mathrm{RV}_{0}$, $L(\lambda)\to 0$, ultimately of bounded variation, and $\theta$ bounded with $\theta(\lambda)\to 0$, ultimately of bounded variation. For $\delta(\lambda)=\lambda L(\lambda)(1+\theta(\lambda))$ one has $\eta(\lambda)=L(\lambda)(1+\theta(\lambda))\in\mathrm{RV}_{0}$ and the conclusions of (c) hold with error $O(L(y/\epsilon))$.
Proof 6.8.
Items (a), (b), (c), (f) follow from Proposition 6.3 with the displayed choices of $\eta\in\mathrm{RV}_{0}$ (uniform convergence for slowly varying functions supplies the replacement $\eta(\Lambda)\rightsquigarrow\eta(y/\epsilon)$). Item (e) follows from Lemma 6.5 with $\ell(\lambda)=\beta(\log(e\lambda))^{\beta-1}$. Item (d) is obtained by a one–step inversion of $y=\epsilon\Lambda-\delta(\Lambda)$ with $\delta=O(1)$, which yields $\Lambda=y/\epsilon+O(1)$, and then by applying Weyl’s law and Karamata’s monotone density theorem to pass to $N_{\mu_{C}}$ and $\rho_{\mathrm{bulk}}$.
Remark 6.9 (Beyond $\mathrm{RV}_{0}$).
Sublinear power terms $\delta(\lambda)=\lambda^{q}$ with $0<q<1$ lie outside the class $\lambda\,\eta(\lambda)$ with $\eta\in\mathrm{RV}_{0}$, but a direct fixed–point estimate gives $\Lambda(C)=\frac{y}{\epsilon}\big(1+O((y/\epsilon)^{q-1})\big),$ hence the same conclusions for $N_{\mu_{C}}$ and $\rho_{\mathrm{bulk}}$ with relative error $O((y/\epsilon)^{q-1})\to 0$.
7 Model geometries and multiplicities
7.1 Constant curvature spaces
Proposition 7.1 (Heat-coefficient alignment for constant curvature).
Let $(M^{d},g)$ be a closed Riemannian manifold of constant sectional curvature $K$ and volume $V:=\mathrm{Vol}(M)$. Its heat trace admits
$\Theta_{\Delta}(t)\sim(4\pi t)^{-d/2}\Big(a_{0}+a_{2}\,t+a_{4}\,t^{2}+\cdots\Big)\qquad(t\downarrow 0),$ (7.1)
with $a_{0}=V$, $a_{2}=\tfrac{1}{6}\!\int_{M}R\,d\mathrm{vol}_{g}=\tfrac{1}{6}d(d-1)K\,V$, and, more generally, $a_{2m}$ a universal polynomial in $K$ times $V$ (odd coefficients vanish for the Laplace–Beltrami operator on a closed manifold). Under the affine encoding $C_{\ell}=\pi-\epsilon\lambda_{\ell}$, the edge-variable heat trace
$H_{\mathrm{edge}}(s)=\sum_{\ell}m_{\ell}e^{-s(\pi-C_{\ell})}=\Theta_{\Delta}(\epsilon s)$ (7.2)
has the small-$s$ expansion
$H_{\mathrm{edge}}(s)\sim(4\pi s)^{-d/2}\Big(\underbrace{\epsilon^{-d/2}a_{0}}_{\text{leading}}+\underbrace{\epsilon^{-(d-2)/2}a_{2}}_{\text{first curvature term}}\,s+\underbrace{\epsilon^{-(d-4)/2}a_{4}}_{\text{quadratic in }K}\,s^{2}+\cdots\Big),\qquad(s\downarrow 0).$ (7.3)
Equivalently, the edge-side coefficients are
$a^{\mathrm{edge}}_{2m}=\epsilon^{\,m-\frac{d}{2}}\,a_{2m},\qquad m=0,1,2,\dots,$ (7.4)
and $a^{\mathrm{edge}}_{2m+1}=0$.
Proof 7.2.
For constant curvature, the Seeley–DeWitt expansion on closed manifolds contains only integer powers; $a_{0}=V$ and $a_{2}=\frac{1}{6}\int R=\frac{1}{6}d(d-1)K\,V$; higher $a_{2m}$ are universal polynomials in the curvature (hence in $K$) times $V$ ([10, 11]). By Proposition 4.4, $H_{\mathrm{edge}}(s)=\Theta_{\Delta}(\epsilon s)$; substituting $t=\epsilon s$ gives
$\Theta_{\Delta}(\epsilon s)=(4\pi\epsilon s)^{-d/2}\sum_{m\geq 0}a_{2m}(\epsilon s)^{m}=(4\pi s)^{-d/2}\sum_{m\geq 0}\epsilon^{\,m-\frac{d}{2}}a_{2m}\,s^{m},$ (7.5)
which yields the stated scaling of coefficients; the vanishing of odd terms persists on the edge side since the manifold has no boundary.
Corollary 7.3 (First two edge-side terms).
With notation as above,
$H_{\mathrm{edge}}(s)\sim(4\pi s)^{-d/2}\Big(\epsilon^{-d/2}\,V+\epsilon^{-(d-2)/2}\,\tfrac{1}{6}d(d-1)K\,V\;s+O(s^{2})\Big),\qquad s\downarrow 0.$ (7.6)
In particular, for $d=2$ one has $H_{\mathrm{edge}}(s)\sim(4\pi s)^{-1}\big(\epsilon^{-1}V+\tfrac{KV}{3}\,s+O(s^{2})\big)$.
Remark 7.4 (Zeta residues).
Using $\zeta_{\mathrm{edge}}(u)=\epsilon^{-u}\zeta_{\Delta}(u)$, the simple poles of $\zeta_{\mathrm{edge}}$ sit at $u=\frac{d-2m}{2}$ with residues $\epsilon^{-(d-2m)/2}\operatorname{Res}_{u=(d-2m)/2}\zeta_{\Delta}(u)$, which are proportional to $a_{2m}$.
7.2 Multiplicity realization via generalized Sturm–Liouville (Kreĭn strings)
Theorem 7.5 (Multiplicity realization via a Kreĭn string).
Let $y_{\ell}:=\pi-C_{\ell}=\epsilon\,\lambda_{\ell}$ and consider the positive discrete measure
$\mu^{+}:=\sum_{\ell\geq 0}m_{\ell}\,\delta_{y_{\ell}}\quad\text{on }(0,\infty),$ (7.7)
where $y_{\ell}\to+\infty$. Then there exists a Kreĭn string (a generalized Sturm–Liouville model with measure coefficients) whose Weyl–Titchmarsh $m$-function is a Stieltjes function having $\mu^{+}$ as its spectral measure. The string is of finite length if and only if $\int(1+t)^{-1}\,d\mu^{+}(t)<\infty$. In particular, the spectrum is pure point and coincides with $\{y_{\ell}\}$ with multiplicities $m_{\ell}$. Under the affine change $C=\pi-y$, the resulting spectral measure equals $\mu_{C}=\sum_{\ell}m_{\ell}\,\delta_{C_{\ell}}$.
Proof 7.6.
Consider the Herglotz–Stieltjes function
$m(z)=\int_{(0,\infty)}\frac{d\mu^{+}(t)}{t-z}=\sum_{\ell\geq 0}\frac{m_{\ell}}{y_{\ell}-z},\qquad z\in\mathbb{C}\setminus(0,\infty).$ (7.8)
By Kreĭn’s correspondence for strings, every Stieltjes function arises as the Weyl–Titchmarsh function of a (finite-length, up to reparametrization) Kreĭn string, and the associated spectral measure is exactly $\mu^{+}$; Kreĭn (1952) and the modern treatment in [13, Thm. 4.1]. Applying the affine change $y=\pi-C$ transports $\mu^{+}$ to $\mu_{C}$.
Theorem 7.7 (Kreĭn uniqueness up to equivalence).
The Kreĭn string realizing $\mu^{+}$ in Theorem 7.5 is unique up to string equivalence (reparametrization of the spatial variable); [13, Thm. 4.1].
Proof 7.8.
Let $m(z)$ denote the Weyl–Titchmarsh function of a (finite-length) Kreĭn string. For a finite string, $m$ is a Stieltjes function and admits an integral representation whose spectral measure is precisely $\mu^{+}$. By [13, Thm. 4.1], the map sending a Kreĭn string (up to equivalence, i.e., reparametrization of the spatial variable) to its Weyl–Titchmarsh function is bijective; in particular, $m(z)$ uniquely determines, and is uniquely determined by, the string up to equivalence. Since $\mu^{+}$ uniquely specifies $m(z)$ (and conversely) via the Stieltjes representation, the realizing Kreĭn string is unique up to equivalence.
Corollary 7.9 (Quadratic encoder sequences are realizable by a Kreĭn string).
Fix $\kappa>0$ and $w_{n}>0$ with $\sum_{n\geq 0}\frac{w_{n}}{1+\kappa(n+1)^{2}}<\infty$ (e.g. $w_{n}\equiv 1$). Let
$x_{n}:=\kappa(n+1)^{2},\qquad\mu^{+}:=\sum_{n\geq 0}w_{n}\,\delta_{x_{n}}.$ (7.9)
Then there exists a (finite-length, up to equivalence) Kreĭn string whose spectral measure equals $\mu^{+}$. Under the affine change $C=\pi-x$, the encoded sequence is
$C_{n}=\pi-x_{n}=\pi-\kappa(n+1)^{2},$ (7.10)
i.e. the quadratic model spectrum $\hat{C}_{n}$ is realized exactly (with weights $w_{n}$).
Proof 7.10.
The integrability $\int(1+t)^{-1}\,d\mu^{+}(t)<\infty$ holds by the displayed summability, which ensures a finite-length string in the sense of [13, Thm. 4.1]. The rest follows from Theorem 7.5 and the affine change $y=\pi-C$.
Lemma 7.11 (No classical smooth SL realization for $d>1$).
Let $\mathcal{L}$ be a classical scalar Sturm–Liouville operator on $[a,b]$ with separated self-adjoint b.c.,
$(p(x)y^{\prime})^{\prime}+q(x)\,y=\lambda\,r(x)\,y,$ (7.11)
where $p,r\in C^{1}([a,b])$, $p>0$, $r>0$, and $q\in L^{1}([a,b])$. Then
$N_{\mathcal{L}}(\Lambda)=\frac{\sqrt{\Lambda}}{\pi}\int_{a}^{b}\!\sqrt{\frac{r(x)}{p(x)}}\,dx\;+\;o(\sqrt{\Lambda})\quad(\Lambda\to\infty),$ (7.12)
hence $\lambda_{n}\sim\big(\frac{\pi n}{\int_{a}^{b}\sqrt{r/p}\,dx}\big)^{2}$. In particular, for $d>1$ no such smooth 1D operator can have an entire spectrum with counting exponent $d/2>1/2$ (equivalently, growth $\lambda_{n}\asymp n^{2/d}$). The same holds for finite matrix SL systems: $N(\Lambda)\sim m\,c\,\sqrt{\Lambda}$.
Proof 7.12.
This is the classical one-dimensional Weyl law; Titchmarsh [16, Ch. XIII] or Levitan–Sargsjan [12]. Under the stated regularity/positivity, the spectrum is discrete and
$N_{\mathcal{L}}(\Lambda)=\frac{1}{\pi}\int_{a}^{b}\!\sqrt{\frac{\Lambda-q(x)}{p(x)}\,r(x)}\,dx\;+\;o(\sqrt{\Lambda})\quad(\Lambda\to\infty).$ (7.13)
The leading term equals $\frac{\sqrt{\Lambda}}{\pi}\int_{a}^{b}\sqrt{r/p}\,dx$, so $N(\Lambda)\sim c\sqrt{\Lambda}$ and $\lambda_{n}\sim c^{\prime}n^{2}$. For a matrix system of size $m$, one only picks up a multiplicative factor $m$ in the leading constant; the exponent remains $1/2$.
Remark 7.13 (Consistency with the companion paper).
The one-dimensional encoder $\hat{C}$ used in the companion paper [1] is a classical smooth Sturm–Liouville operator on a compact interval with eigenvalues $C_{n}=\pi-\kappa(n+1)^{2}$, decreasing quadratically in $n$. This does not reproduce the full encoded spectrum $\{C_{\ell}\}$ for $d>1$, where $\pi-C_{\ell}\sim\kappa\,\ell^{2/d}$ decreases more slowly (since $2/d<2$). The role of $\hat{C}$ is to provide an affine spectral encoding: after the pushforward $C=\pi-\epsilon\lambda$, the bulk edge-variable Weyl scaling $N_{\mu_{C}}(C)\sim\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2}$ (and the density exponent (d2)/2(d-2)/2) follows from the original Weyl law, independently of low-lying modes near $C\uparrow\pi$. An exact realization of the entire discrete measure $\mu_{C}$ with multiplicities is instead furnished by the generalized one-dimensional model (Kreĭn string) in Theorem 7.5.
7.3 Spectral clustering remarks
On the round sphere $S^{d}$ the Laplace eigenvalues form clusters $\lambda_{\ell}=\ell(\ell+d-1)$ with multiplicities [2, Ch. 1]
$m_{\ell}=\binom{\ell+d}{\ell}-\binom{\ell+d-2}{\ell-2}=\frac{2\ell+d-1}{\ell}\binom{\ell+d-2}{d-1}\quad(\ell\geq 1),\qquad m_{0}=1.$ (7.14)
Under the affine encoding $C_{\ell}=\pi-\epsilon\,\lambda_{\ell}$ the spacing in the $C$-variable is
$C_{\ell}-C_{\ell+1}=\epsilon\big(\lambda_{\ell+1}-\lambda_{\ell}\big)=\epsilon\,(2\ell+d),$ (7.15)
so steps of the counting function $N_{\mu_{C}}$ occur at an increasingly sparse lattice as $\ell$ grows, while the heights of the steps (the multiplicities) increase polynomially (indeed $m_{\ell}\sim\frac{2}{(d-1)!}\,\ell^{d-1}$).
Lemma 7.14 (Average multiplicity in a short $C$-window).
Let $\delta=\delta(C)>0$. Assume $\delta(C)=o(\pi-C)$ as $C\uparrow\pi$ and $\delta(C)=o(|C|^{1/2})$ as $C\to$-$\infty$. Define, whenever the window contains at least one cluster,
$\overline{m}(C,\delta):=\frac{N_{\mu_{C}}(C)-N_{\mu_{C}}(C-\delta)}{\#\{\ell:\;C_{\ell}\in[C-\delta,C]\}}.$ (7.16)
Then:
  1. (i)
    As $C\uparrow\pi$, one has $\{\ell:\,C_{\ell}\in[C-\delta,C]\}=O(1)$ and $\overline{m}(C,\delta)=O(1)$. In particular, if $\delta<\epsilon d$, the window contains at most one cluster and hence $\overline{m}(C,\delta)=m_{\ell_{\ast}}$ for some fixed $\ell_{\ast}\in\{0,1,\dots\}$ (e.g. $m_{0}=1,\,m_{1}=d+1$).
  2. (ii)
    As $C\to$-$\infty$, let $\ell(C)$ be the (unique) index with $\lambda_{\ell(C)}\leq(\pi-C)/\epsilon<\lambda_{\ell(C)+1}$. Then
    $\overline{m}(C,\delta)\ \sim\ m_{\ell(C)}\ \sim\ \frac{2}{(d-1)!}\Big(\frac{|C|}{\epsilon}\Big)^{\frac{d-1}{2}}.$ (7.17)
Proof 7.15.
Write $\Lambda:=(\pi-C)/\epsilon$ and $\Delta:=\delta/\epsilon$. Since $C=\pi-\epsilon\lambda$ is strictly decreasing in $\lambda$, the set $\{\ell:\,C_{\ell}\in[C-\delta,C]\}$ equals
$\{\ell:\,\lambda_{\ell}\in[\Lambda,\ \Lambda+\Delta]\},$ (7.18)
up to the choice of open/closed endpoints (which does not affect counting with multiplicities). Denote this set by $I(C,\delta)$ and its cardinality by $K(C,\delta)$.
Step 1: upper bounds for $K(C,\delta)$. Let $s_{\min}:=\min_{\ell\in I(C,\delta)}(\lambda_{\ell+1}-\lambda_{\ell})$. Then any interval in the $\lambda$-axis of length $\Delta$ contains at most $1+\Delta/s_{\min}$ grid points $\{\lambda_{\ell}\}$, hence
$K(C,\delta)\ \leq\ 1+\frac{\Delta}{s_{\min}}.$ (7.19)
Near the edge $C\uparrow\pi$ we have $\Lambda\downarrow 0$, so $I(C,\delta)\subset\{0,1,\dots,L_{0}\}$ with some fixed $L_{0}$ for all $C$ close to $\pi$; in this band $s_{\min}\geq d$. Since $\delta=o(\pi-C)$ implies $\Delta\to 0$, (7.19) gives $K(C,\delta)=O(1)$. Moreover, if $\delta<\epsilon d$ then $\Delta<d\leq s_{\min}$ and $K(C,\delta)\leq 1$.
As $C\to$-$\infty$, the relevant indices satisfy $\ell\asymp\sqrt{\Lambda}$. Hence throughout $I(C,\delta)$ we have $s_{\min}\asymp 2\ell\asymp 2\sqrt{\Lambda}$. From (7.19) and $\Delta=\delta/\epsilon$ we get
$K(C,\delta)\ \leq\ 1+\frac{\delta/\epsilon}{c\sqrt{\Lambda}}\ \sim\ 1+\frac{\delta}{c\sqrt{\epsilon|C|}},$ (7.20)
which is $O(1)$ and in fact equals $1$ for all sufficiently negative $C$ if $\delta=o(|C|^{1/2})$. In particular, the window eventually contains at most one cluster.
Step 2: control of multiplicities in the window. Near the edge $C\uparrow\pi$ the set of admissible indices $I(C,\delta)$ is contained in a fixed finite set; hence $\sup_{\ell\in I(C,\delta)}m_{\ell}\leq M_{0}$ for some constant $M_{0}$, which yields $\overline{m}(C,\delta)=O(1)$. If furthermore $K(C,\delta)=1$, then trivially $\overline{m}(C,\delta)=m_{\ell_{\ast}}$ for the unique $\ell_{\ast}\in I(C,\delta)$.
For $C\to$-$\infty$, we use the asymptotic $m_{\ell}\sim\frac{2}{(d-1)!}\ell^{d-1}$. Since $K(C,\delta)$ is bounded and, under $\delta=o(|C|^{1/2})$, eventually equals $1$, the indices in $I(C,\delta)$ form a set of the form $\{\ell(C)+j:\,j\in J\}$ with $J\subset\mathbb{Z}$ finite and independent of $C$. Therefore,
$\frac{m_{\ell(C)+j}}{m_{\ell(C)}}\ \longrightarrow\ 1\qquad(C\to-\infty)$ (7.21)
for each fixed $j$, because $m_{\ell}$ is a polynomial in $\ell$ of degree d1d-1. It follows that
$\overline{m}(C,\delta)\ =\ \frac{1}{K(C,\delta)}\sum_{\ell\in I(C,\delta)}m_{\ell}\ =\ m_{\ell(C)}\cdot\frac{1}{K(C,\delta)}\sum_{j\in J}\frac{m_{\ell(C)+j}}{m_{\ell(C)}}\ \longrightarrow\ m_{\ell(C)}.$ (7.22)
Finally, $\ell(C)\sim\sqrt{\Lambda}\sim\sqrt{|C|/\epsilon}$, hence $m_{\ell(C)}\sim\frac{2}{(d-1)!}\big(\frac{|C|}{\epsilon}\big)^{\frac{d-1}{2}}$, which yields the stated asymptotics.
Thus the microscopic staircase due to clustering is negligible near the geometric edge ($C\uparrow\pi$), where jumps are small and sparse, and it averages out in the bulk ($C\to$-$\infty$), which is favorable for practical extraction of $d$ and $\gamma_{d}$ from bulk edge-variable data (cf. Corollary 3.3). In the Kreĭn string realization (Theorem 7.5), multiplicities are encoded as the weights of the atomic spectral measure, so clustering is faithfully carried by $\mu_{C}$ without affecting the leading asymptotics.
Proposition 7.16 (Edge hit probability for a fixed cluster).
Let $C_{\ell}=\pi-\epsilon\,\lambda_{\ell}$ with $\lambda_{\ell}=\ell(\ell+d-1)$, and let $\delta=\delta(C)>0$. Fix an index $\ell$ and consider the following randomization: choose $U$ uniformly on the interval $0,\,C_{\ell}-C_{\ell+1})$ and set the window $C-\delta,\,C]$ with $C=C_{\ell}-U$. If $\delta<C_{\ell}-C_{\ell+1}=\epsilon(2\ell+d)$, then
$\mathbb{P}\big(\,C_{\ell}\in[C-\delta,\,C]\,\big)=\frac{\delta}{C_{\ell}-C_{\ell+1}}=\frac{\delta}{\epsilon(2\ell+d)}.$ (7.23)
In particular, for fixed (or uniformly bounded) $\delta$, this probability decays like $O(\ell^{-1})$ as $\ell\to\infty$.
Proof 7.17.
Under the stated randomization, the event $\{C_{\ell}\in[C-\delta,C]\}$ is equivalent to $\{U\in[0,\delta]\}$ inside a fundamental cell of length $C_{\ell}-C_{\ell+1}$. Since $U$ is uniform, the probability equals $\delta/(C_{\ell}-C_{\ell+1})$. The identity $C_{\ell}-C_{\ell+1}=\epsilon(2\ell+d)$ follows from $\lambda_{\ell+1}-\lambda_{\ell}=2\ell+d$. The large-$\ell$ decay is immediate.
Remark 7.18 (Jump-size weighting inside a window).
If one conditions on the event that the window contains exactly one cluster and then samples a jump of $N_{\mu_{C}}$ uniformly among the atoms inside the window, the observed jump size equals the multiplicity $m_{\ell}$ of that cluster. Thus, conditional on “one-cluster” windows, the distribution of observed jump sizes is the empirical distribution of $\{m_{\ell}\}$ at the corresponding indices, not $1/m_{\ell}$. Near the edge $C\uparrow\pi$, small windows contain only $\ell=0$ (and then $\ell=1$), so the observed jump is $m_{0}=1$ (then $m_{1}=d+1$), in agreement with Lemma 7.14.
8 Examples
8.1 The round sphere $S^{d}$
Corollary 8.1 (Explicit $\gamma_{d}$ on $S^{d}$).
For the unit round sphere $S^{d}$,
$\gamma_{d}=\frac{\omega_{d}}{(2\pi)^{d}}\,\mathrm{Vol}(S^{d})=\frac{2^{\,1-d}\,\sqrt{\pi}}{\Gamma\!\big(\tfrac{d}{2}+1\big)\,\Gamma\!\big(\tfrac{d+1}{2}\big)}$ (8.1)
[2, 6]. In particular, for $d=3$ one has $\gamma_{3}=\tfrac{1}{3}$. Hence, as $C\to$-$\infty$,
$N_{\mu_{C}}(C)\sim\tfrac{1}{3}\,\epsilon^{-3/2}(\pi-C)^{3/2},\qquad\rho_{\mathrm{bulk}}(C)\sim\tfrac{1}{2}\,\epsilon^{-3/2}(\pi-C)^{1/2}\quad(d=3).$ (8.2)
Proof 8.2.
Use $\omega_{d}=\pi^{d/2}/\Gamma(\tfrac{d}{2}+1)$ and $\mathrm{Vol}(S^{d})=2\pi^{(d+1)/2}/\Gamma(\tfrac{d+1}{2})$, then substitute into $\gamma_{d}=\frac{\omega_{d}}{(2\pi)^{d}}\mathrm{Vol}(S^{d})$. For $d=3$ this gives $\gamma_{3}=1/3$. Insert into Theorem 3.1.
Corollary 8.3 (Numerical sanity check for $S^{3}$).
For the unit sphere $S^{3}$ one has $\gamma_{3}=\tfrac{1}{3}$, $\lambda_{\ell}=\ell(\ell+2)$ and $m_{\ell}=(\ell+1)^{2}$. Fix $\epsilon=1$ and set $y=\pi-C$, and, for illustration, take $\delta=0.1\,y$, so that the $\lambda$–window is $\Lambda,\Lambda+\delta]$ with $\Lambda=y$.
  1. (a)
    Edge regime: if $y=0.1$ then $\Lambda,\Lambda+\delta]=[0.1,0.11]$ contains no eigenvalues ($\lambda_{0}=0$, $\lambda_{1}=3$), hence the window produces no jump.
  2. (b)
    Away from the edge: if $y=8$ then $\Lambda,\Lambda+\delta]=[8,8.8]$ contains $\lambda_{2}=8$, so the unique jump has size $m_{2}=(2+1)^{2}=9$ and the observed step is $9$.
Both observations are consistent with Lemma 7.14.
Remark 8.4 (3D bulk case).
For $d=3$, Corollary 8.1 gives $N_{\mu_{C}}(C)\sim\tfrac{1}{3}\,\epsilon^{-3/2}(\pi-C)^{3/2}$ and $\rho_{\mathrm{bulk}}(C)\sim\tfrac{1}{2}\,\epsilon^{-3/2}(\pi-C)^{1/2}$. Thus the bulk density exponent $1/2$ recovers $d=3$ from edge-variable data via Theorem 3.1. The quadratic encoder model $C_{n}=\pi-\kappa(n+1)^{2}$ can be realized by a Kreĭn string (Corollary 7.9); the affine pushforward $C=\pi-\epsilon\lambda$ transfers the geometric Weyl scaling independently of low modes near $C\uparrow\pi$.
8.2 The flat torus $T^{d}$
Corollary 8.5 (Flat torus $T^{d}$).
For a flat torus $T^{d}=\mathbb{R}^{d}/\Lambda$ with volume $\mathrm{Vol}(T^{d})$ one has $\gamma_{d}=\frac{\omega_{d}}{(2\pi)^{d}}\,\mathrm{Vol}(T^{d})$. Consequently, as $C\to$-$\infty$,
$N_{\mu_{C}}(C)\sim\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2},\qquad\rho_{\mathrm{bulk}}(C)\sim\tfrac{d}{2}\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{\frac{d-2}{2}}.$ (8.3)
Here $\omega_{d}=\mathrm{Vol}(B_{d})=\pi^{d/2}/\Gamma(\tfrac{d}{2}+1)$.
Proof 8.6.
Use the standard Weyl law for flat tori [2] and Theorem 3.1.
8.3 Compact hyperbolic surfaces
Let $X$ be a compact hyperbolic surface … Weyl’s law in dimension 22 reads $N_{\Delta}(\Lambda)\sim\frac{\mathrm{Area}(X)}{4\pi}\,\Lambda$ [14, 15]. Hence, as $C\to$-$\infty$,
$N_{\mu_{C}}(C)\sim\frac{\mathrm{Area}(X)}{4\pi}\,\epsilon^{-1}\,(\pi-C),\qquad\rho_{\mathrm{bulk}}(C)\sim\frac{\mathrm{Area}(X)}{4\pi}\,\epsilon^{-1}.$ (8.4)
Thus the bulk density is asymptotically constant (exponent (d2)/2=0(d-2)/2=0 for $d=2$), independently of irregular multiplicity patterns arising from symmetries, in agreement with Theorem 3.1.
8.4 Berger spheres
Corollary 8.7 (Bulk Weyl asymptotics on Berger spheres).
Let $S^{3}_{k}$ denote the Berger sphere with deformation parameter $k>0$, equipped with the left-invariant metric
$ds^{2}=\sigma_{1}^{2}+\sigma_{2}^{2}+k^{2}\sigma_{3}^{2},$ (8.5)
where $\{\sigma_{i}\}$ are the standard left-invariant one-forms on $SU(2)$ [17]. The Laplace spectrum on functions is given by
$\lambda_{n,m}(k)=n(n+2)+(k^{2}-1)m^{2},\qquad|m|\leq n,$ (8.6)
with multiplicity $n+1$. As $C\to$-$\infty$, one recovers the universal bulk asymptotics
$N_{\mu_{C}}(C)\ \sim\ \tfrac{1}{3}\,\epsilon^{-3/2}(\pi-C)^{3/2},\qquad\rho_{\mathrm{bulk}}(C)\ \sim\ \tfrac{1}{2}\,\epsilon^{-3/2}(\pi-C)^{1/2}.$ (8.7)
Remark 8.8.
The deformation parameter $k$ affects low-lying eigenvalues but does not alter the bulk scaling, in agreement with Theorem 3.1. This illustrates robustness of the affine encoding under smooth metric deformations. For visualization, compare log-log plots of $\rho_{\mathrm{bulk}}(C)$ for $k=1$ (standard $S^{3}$) and $k\neq 1$, showing identical bulk exponents.
8.5 Lens spaces
Corollary 8.9 (Bulk Weyl asymptotics on lens spaces).
Let $L(p,q)=S^{3}/\mathbb{Z}_{p}$ be a 3-dimensional lens space with the induced round metric [18]. Its Laplace spectrum is a subset of that of $S^{3}$ with reduced multiplicities, and
$\mathrm{Vol}(L(p,q))=\tfrac{1}{p}\,\mathrm{Vol}(S^{3}).$ (8.8)
Hence, as $C\to$-$\infty$,
$N_{\mu_{C}}(C)\ \sim\ \tfrac{1}{3p}\,\epsilon^{-3/2}(\pi-C)^{3/2},\qquad\rho_{\mathrm{bulk}}(C)\ \sim\ \tfrac{1}{2p}\,\epsilon^{-3/2}(\pi-C)^{1/2}.$ (8.9)
Remark 8.10.
Although different lens spaces $L(p,q)$ may be non-isometric but isospectral, the edge-variable asymptotics depend only on $d$ and the volume. This confirms that affine encoding preserves classical Weyl invariants while ignoring finer topological features. For example, compare $p=1$ (standard $S^{3}$) with $p>1$, where the bulk density scales inversely with $p$, reflecting the topological quotient.
8.6 Manifolds with boundary: the ball
Proposition 8.11 (Dirichlet Laplacian on the ball $B^{3}$).
Consider the unit ball $B^{3}\subset\mathbb{R}^{3}$ with Dirichlet boundary conditions [8]. Weyl’s two-term law gives
$N_{\Delta}(\Lambda)=\tfrac{\omega_{3}}{(2\pi)^{3}}\,\mathrm{Vol}(B^{3})\,\Lambda^{3/2}-\tfrac{\omega_{2}}{4(2\pi)^{2}}\,\mathrm{Area}(S^{2})\,\Lambda+o(\Lambda).$ (8.10)
Under affine encoding $C=\pi-\epsilon\lambda$ this transfers to
$N_{\mu_{C}}(C)\ \sim\ \tfrac{1}{3}\,\epsilon^{-3/2}(\pi-C)^{3/2}-\tfrac{1}{16}\,\epsilon^{-1}(\pi-C)+o(\pi-C).$ (8.11)
Remark 8.12.
The boundary term survives the transfer and appears as the linear correction in the bulk expansion, confirming Proposition 3.4. This example highlights how the method handles manifolds with boundary, with the edge-variable asymptotics capturing both volume and area contributions. A numerical plot of $N_{\mu_{C}}(C)$ vs. $(\pi-C)$ would visually separate the bulk and boundary terms.
9 Conclusion
The analysis establishes that the edge-variable encoding $C=\pi-\epsilon\lambda$ reproduces Weyl asymptotics with coefficient $\gamma_{d}$ and bulk density exponent (d2)/2(d-2)/2, that this scaling is stable under admissible perturbations, and that for constant curvature spaces the heat coefficients transfer exactly to the edge-side expansion. Multiplicity patterns, although not realizable by smooth Sturm–Liouville operators in dimensions $d>1$, admit a complete representation by Kreĭn strings unique up to equivalence; clustering on spheres averages out in the bulk and does not alter the limiting law. Explicit evaluations of $\gamma_{d}$ for spheres, tori, hyperbolic surfaces, as well as deformed and factored geometries (Berger spheres, lens spaces) and bounded domains, confirm consistency with the reconstruction formula. Open problems include the full extension to general manifolds with boundary, operators beyond the Laplacian, and finer inverse spectral questions.